# Finch members

### v3

@@ 4 Related Work
The integration of LLMs into enterprise productivity tools has accelerated dramatically recently. The recently launched ChatGPT Agent [openai_agent] extends these capabilities to autonomous task completion, enabling multi-step workflows across web browsing, code execution, and spreadsheet manipulation. Microsoft Copilot [microsoft_copilot] embeds AI capabilities across the Microsoft 365 suite, enabling users to draft documents, analyze spreadsheets, and automate workflows through natural language interaction. Similarly, Google has integrated Gemini [gemini_workspace] into Google Workspace, providing AI-assisted features in Docs, Sheets, and Gmail. Anthropic’s Claude Excel has also entered the enterprise space with spreadsheet automation capabilities [claude_excel], while remarkable tools like Shortcut AI [shortcut_spreadsheet] focus specifically on AI-powered spreadsheet manipulation.
The emergence of agentic AI systems marks a significant shift from understanding and QA to autonomous task completion [li2023sheetcopilot, chen2024sheetagent, ma2024spreadsheetbench]. However, there are long-standing challenges such as messy inputs and multimodal processing. SpreadsheetLLM [spreadsheetllm2024] introduces novel encoding and compression methods to help LLMs understand large and messy spreadsheet structures, and further addresses this challenge through spreadsheet post-training. Beyond structural understanding, multimodal processing remains challenging for spreadsheet AI systems. On the formatting front, early work explored neural approaches for table formatting [dong2020neural]. Recent advances in formula generation have progressed from pretraining with numerical reasoning [cheng2022fortap] to natural language-driven formula synthesis [zhao2024nl2formula], contrastive learning-based recommendation [chen2024auto], and interactive formula prediction through hierarchical expansion [he2023hermes].
Recent years have seen significant progress in benchmarks for financial reasoning [chen2021finqa, islam2023financebench, finance_agent_benchmark, finben2024, findabench2025, finagentbench2025, finauditing2025, li2025investorbench, visfineval2025, gdpval2025, finsearchcomp2025, xfinbench2025, xie2024finben], spreadsheet reasoning [li2023sheetcopilot, chen2024sheetagent, ma2024spreadsheetbench, sodbench2025, li2023auto, dong2019tablesense, sheetmind2025, spreadsheetllm2024, wu2025realhitbench, li2025mimotable, dong2025machinelearninglm, dong2025reasoning, li2024tapilot, zhao2024nl2formula, odysseybench2025, officebench2024], and multimodal document [mathew2021docvqa], table [li2020tablebank, zheng2024multimodal], chart [wang2024charxiv], table-chart [dong2024ttc], and spreadsheet [xia2024vision, dong2019tablesense] reasoning, driving advances in LLM-based agents for enterprise tasks. However, Finch proposes a new benchmark for the messy artifacts and long-horizon workflows in wild F&A enterprise settings.
@@ 5 Conclusion
We introduced Finch, a new benchmark for real-world F&A enterprise workflows. Finch combines workflows induced from enterprise email threads, version histories of spreadsheets, and high-quality financial artifacts with rigorous expert annotation and a calibrated LLM-as-judge framework, enabling systematic evaluation of agents on diverse workflows that operate over large, messy, and multimodal enterprise artifacts and require long-horizon, spreadsheet-centric reasoning. Our experiments show that even the strongest frontier systems pass fewer than 40% of workflows after spending 16.8 minutes per workflow, revealing a substantial gap between current AI capabilities and the demands of real enterprise practice. We hope Finch will serve as a foundation for developing agents to tackle real, messy and long-horizon professional work.
@@ Appendix A Author List
*Corresponding author.
@@ Appendix C Detailed Analysis
Appendix D Ethics Statement
@@
Appendix E Examples
E.1 Example 1

### v4

@@ 4 Related Work
The integration of LLMs into enterprise productivity tools has accelerated dramatically recently. The recently launched ChatGPT Agent [41] extends these capabilities to autonomous task completion, enabling multi-step workflows across web browsing, code execution, and spreadsheet manipulation. Microsoft Copilot [39] embeds AI capabilities across the Microsoft 365 suite, enabling users to draft documents, analyze spreadsheets, and automate workflows through natural language interaction. Similarly, Google has integrated Gemini [22] into Google Workspace, providing AI-assisted features in Docs, Sheets, and Gmail. Anthropic’s Claude Excel has also entered the enterprise space with spreadsheet automation capabilities [2].
The emergence of agentic AI systems marks a significant shift from understanding and QA to autonomous task completion [30, 7, 37]. However, there are long-standing challenges such as messy inputs and multimodal processing. SpreadsheetLLM [17] introduces novel encoding and compression methods to help LLMs understand large and messy spreadsheet structures, and further addresses this challenge through spreadsheet post-training. Beyond structural understanding, multimodal processing remains challenging for spreadsheet AI systems. On the formatting front, early work explored neural approaches for table formatting [15]. Recent advances in formula generation have progressed from pretraining with numerical reasoning [9] to natural language-driven formula synthesis [57], contrastive learning-based recommendation [6], and interactive formula prediction through hierarchical expansion [23].
Recent years have seen significant progress in benchmarks for financial reasoning [8, 27, 5, 53, 35, 10, 46, 29, 36, 43, 25, 56, 54], spreadsheet reasoning [30, 7, 37, 26, 33, 13, 59, 17, 50, 34, 16, 12, 31, 57, 45, 47], and multimodal document [38], table [32, 58], chart [48], table-chart [14], and spreadsheet [52, 13] reasoning, driving advances in LLM-based agents for enterprise tasks. However, Finch targets messy artifacts and long-horizon workflows in real-world finance and accounting settings.
@@ 5 Conclusion
We introduced Finch, a new benchmark for real-world F&A enterprise workflows. Finch combines workflows induced from enterprise email threads, version histories of spreadsheets, and high-quality financial artifacts with rigorous expert annotation and a calibrated LLM-as-judge framework, enabling systematic evaluation of agents on diverse workflows that operate over large, messy, and multimodal enterprise artifacts and require long-horizon, spreadsheet-centric reasoning. Our experiments show that even the strongest frontier systems pass fewer than 50% of workflows after spending 16.8 minutes per workflow, revealing a substantial gap between current AI capabilities and the demands of real enterprise practice. We hope Finch will serve as a foundation for developing agents to tackle real, messy and long-horizon professional work.
6 Acknowledgments
We thank Chi Zhang and Jieqiong Li for reviewing the dataset and providing valuable advice.
@@ Appendix A Author List
Author Affiliation Email
Haoyu Dong* UCAS; Microsoft donghaoyu82@gmail.com
Pengkun Zhang South China University of Technology sezhangpengkun@mail.scut.edu.cn
Yan Gao Zhongguancun Academy gaoyan@zgci.ac.cn
Xuanyu Dong Harvest Fund qianmuxuanyu@126.com
Yilin Cheng Fudan University; Zhongguancun Academy ylcheng23@m.fudan.edu.cn
Mingzhe Lu UCAS lumingzhe23@mails.ucas.edu.cn
Zikun Zhu Duke University zz295@duke.edu
Adina Yakefu Hugging Face adina@huggingface.co
Shuxin Zheng Zhongguancun Academy sz@bjzgca.edu.cn
*Corresponding author. UCAS stands for the University of Chinese Academy of Sciences
@@ Appendix C Detailed Analysis
Appendix D Tool-Call Overhead Case Study
To better understand the practical complexity of completing Finch tasks, we conduct a small-scale case study on tool-call overhead. From the full set of 172 workflows, we select 20 representative examples covering different task types and varying levels of complexity, and analyze how many tool calls are required for an LLM-based agent to complete them.
D.1 Methodology
For each selected task, we use the up-to-date (March 2026) Claude Coworker (Opus 4.6) in an isolated session and ask it to solve the task according to the provided instruction. The agent is also required to write its tool-use trajectory into a log file, recording each tool call step by step. Since the number of tool calls in a failed attempt may differ from that in a successful completion, we use a second session to compare the produced output against the reference answer file and determine whether the result is correct.
When the output is incorrect, we observe two distinct cases from the perspective of tool usage. In the first case, the overall tool-use procedure is already complete, but the execution within one or more steps is wrong; for example, the agent invokes a Python script as needed, but the script contains an incorrect formula. In the second case, the trajectory is missing one or more critical steps that are necessary for producing the correct output. This second case is relatively uncommon, appearing only twice among the 20 examples we analyze. For such cases, we ask the model to identify the missing indispensable steps and merge them into the original trajectory. Therefore, our final count is intended to approximate the tool calls required to correctly complete the task, rather than merely reporting the raw tool calls from a failed run.
D.2 Results and Analysis
Table 6 summarizes the 20 selected tasks and their corresponding numbers of tool calls.
Workflow ID Task Type(s) #Tasks #Tool Calls
46 Calculation 1 7
31 Summary / Visualization 1 8
26 Summary / Visualization 1 10
47 Calculation 1 13
17 Cross-sheet/file Retrieval 1 14
137 Structuring / Formatting 1 16
91 Calculation, Financial Modeling 2 6
106 Structuring / Formatting, Calculation 2 6
49 Validation / Review, Calculation 2 14
16 Calculation, Summary / Visualization 2 17
28 Validation / Review, Structuring / Formatting 2 17
158 Validation / Review, Calculation 2 18
15 Structuring, Translation, Summary / Vis. 3 9
150 Calculation, Data Entry, Cross-sheet Retrieval 3 9
66 Calculation, Structuring, Financial Modeling 3 17
139 Financial Modeling, Calculation, Summary / Vis. 3 25
159 Structuring, Calc., Fin. Modeling, Validation 4 14
81 Data Entry, Structuring, Validation, Calc. 4 17
19 Web Search, Data Entry, Structuring, Calc. 4 107
88 Web Search, Data Entry, Structuring, Calc., Vis. 5 71
Table 6: Case study of tool-call overhead on 20 representative tasks.
Web-grounded tasks incur substantially higher tool-call overhead.
Among the 20 cases, the two tasks involving web search (ID 19 and ID 88) require dramatically more tool calls than the others. In Task 19, websearch and webfetch together account for 60 out of 107 tool calls (56%). In Task 88, these two tools account for 34 out of 71 tool calls (48%). During inspection, we find that web-grounded tasks often require repeated retrieval from multiple external sources, followed by cross-source comparison, integration, and verification. As a result, the dominant cost in such tasks comes not from spreadsheet manipulation itself, but from external information acquisition and validation.
Distribution of tool calls for non-web tasks.
After excluding the two web-search tasks, the number of tool calls ranges from 6 to 25, with a mean of 13.2 and a median of 14. The first and third quartiles are 9 and 17, respectively, giving an interquartile range of 8.
Task-type count alone does not explain tool-call overhead.
After removing the two web-search outliers, we do not observe a clear positive relationship between the number of task-type labels and the number of tool calls. For example, both Task 46 and Task 47 are labeled as Calculation, but Task 47 requires nearly twice as many tool calls due to additional business-logic reasoning (sign conventions, Purchase vs. Pooling Method tax-base logic, and cross-checking of business consistency).
Limitations.
This analysis should be viewed as a case study rather than a full-scale statistical evaluation. The 20 tasks are selected as representative examples rather than a random sample, and the measured tool-call counts depend on the behavior of a specific agent configuration (Claude Coworker, Opus 4.6). In addition, our corrected trajectories for failed runs involve a post hoc reconstruction of missing indispensable steps, which introduces some degree of judgment into the counting process.
D.3 Representative Tool-Call Logs
We present two representative tool-call logs. Task 26 illustrates a standard successful trajectory; Task 139 illustrates a case where the original trajectory omits critical steps (marked with + Bash) that must be added for correct completion.
D.3.1 Task 26: Chart Generation across Scenario Sheets
Task.
Add two charts to each of the three worksheets (Scenario1--3): (1) a Hedge Coverage stacked column chart (first year; X = months; stacked series = contract MW; Short = reference line; Y = MW), and (2) a Portfolio Cost line chart (all years; X = months; series = Portfolio Weighted Avg and Monthly Price; Y = \$/MWh).
Tool-call log  (10 calls) 1. Bashls mnt/26/ — confirm source file exists. 2. Read metadata.json — parse task instructions and output requirements. 3. Read SKILL.md — review best practices for Excel operations. 4. Bash Python: inspect workbook structure via openpyxl (sheet names, first 20 rows, column mapping). 5. Bash Read full output (previous step was truncated). 6. Bash Python: inspect first 20 rows of each Scenario sheet in detail—column semantics, data types. 7. Bash Python: read date column (Z, rows 4–63) to identify first-year range (rows 4–15) and full range (Jan 2001–Dec 2005). 8. Bash cp 26_src_0.xlsx 26.xlsx — preserve original, work on copy. 9. Bash Run add_charts.py: create all 6 charts (2 per Scenario sheet) via openpyxl. 10. Bash Python: verify chart count, titles, types, and series per sheet.
D.3.2 Task 139: XNPV Sensitivity Analysis with Corrected Trajectory
Note.
This is a corrected tool-call log. Steps marked + Bash were missing from the original trajectory and added post hoc; they mainly involve constructing monthly cash-flow data for XNPV Variables, building multi-year analysis data for Cedar Brakes, and writing native Excel XNPV formulas.
Tool-call log  (25 calls)  Phase 1 — Understand the workbook 1. Bashls mnt/139/ — confirm source file. 2. Read metadata.json — task instructions, source file list, output requirements. 3. Read SKILL.md — Excel operation best practices. 4. Bash Python (data_only=True): inspect all sheet names, dimensions, and data values. 5. Bash Python: read Damage Calculations sheet—annual capacity payments (Q1–Q4), XNPV values, sensitivity-table region. 6. Bash Python: read Summary sheet—layout of Q1–Q4 sections and Delta region. 7. Bash Python (data_only=False): read formulas in Damage Calculations—calculation logic, DataTableFormula objects. 8. Read Read full output of step 7 (truncated). 9. Bash Python: read formulas in Summary sheet—formula reference relationships. 10. Bash Python: read formulas and variables in Tenaska’s i Adjustment sheet. 11. Bash Python: read variable values in Tenaska’s i Adjustment. 12. Bash Python: read formulas and variables in Brazos’ i Adjustment—compare with Tenaska. 13. Bash Python: validate XNPV calculation method—date conventions, discounting logic. 14. Bash Python: inspect DataTableFormula parameters—confirm input variable is F8 (plant capacity).  Phase 2 — Build missing intermediate data (added steps) 15. + BashPython: read Cedar Brakes sheet—annual data layout (Year, Volume, Old PPA, Market, SI, New PPA, Savings), column relationships, year range (2000–2016+), and rows/columns to fill. 16. + Bash Python: compute annual capacity payments for Q1–Q4 under four capacity assumptions (240/245/258/263 MW), applying the rate structure and interest-rate adjustments from Tenaska and Brazos. 17. + Bash Python: convert annual payments into monthly cash-flow sequences (date sequence from Jan 1997 to contract end; monthly = annual/12); verify totals.  Phase 3 — Execute main script 18. TodoWriteCreate task list. 19. BashMain Python script—copy source to 139.xlsx, then: + XNPV Variables Write title, headers (Date, Q1--Q4), and monthly cash-flow rows starting at row 88 of Damage Calculations; scale payments proportionally for 240/245/258/263 MW. + Cedar Brakes Fill multi-year financial data (rows 24+): Volume, Old/New PPA, Market, SI, Savings for all years 2000–end. + XNPV formulas Write =XNPV(rate, flows, dates) formulas into the sensitivity-table region, referencing the cash-flow data above. Original Fill Q1–Q4 capacity values and Delta section (differences vs. 263 MW baseline) in Summary; save. 20. TodoWriteUpdate progress. 21. Glob Locate recalc.py. 22. Bash python recalc.py — recalculate all Excel formulas (including XNPV).  Phase 4 — Verify outputs 23. + BashPython (data_only): verify XNPV Variables integrity (row count, date continuity, cash-flow values), Cedar Brakes data, and XNPV formula results. 24. Bash Python: verify Summary (PV5, debt, SWAP, implied equity for Q1–Q4) and Damage Calculations sensitivity table (all XNPV values, Delta vs. 263 MW baseline). 25. TodoWrite Final progress update.
Appendix E Ethics Statement
@@
Appendix F Examples
This section presents representative examples from Finch to illustrate the diversity and complexity of spreadsheet-centric enterprise workflows covered by the benchmark. The examples span a wide range of realistic professional tasks, including cross-sheet verification, formula auditing, document-grounded extraction, schema transformation, financial modeling, and reporting.
Note that the figures shown in this section are illustrative excerpts rather than complete workflow inputs or full task specifications. Many Finch workflows involve large, multi-sheet or multi-file inputs and extensive supporting documents, which cannot be fully visualized in static screenshots. Accordingly, the figures are intended to highlight representative data characteristics, structural complexity, and reasoning challenges that arise in the underlying workflows, rather than to fully specify the tasks themselves.
Across these workflows, AI agents are required to jointly reason over heterogeneous artifacts, while preserving structural consistency and semantic correctness. Many tasks demand multi-step reasoning, precise interpretation of formulas, and careful cross-referencing across sheets or external documents. Collectively, these examples highlight the core challenges targeted by Finch: messy and multimodal document processing, long-horizon reasoning, robust code generation under schema variation, and faithful execution of complex financial and accounting logic.
F.1 Example 1